Advertisements
Advertisements
प्रश्न
Suppose the chances of hitting a target by a person X is 3 times in 4 shots, by Y is 4 times in 5 shots, and by Z is 2 times in 3 shots. They fire simultaneously exactly one time. What is the probability that the target is damaged by exactly 2 hits?
Advertisements
उत्तर
Given
Probability X hitting the target P(X) = `3/4`
Probability Y hitting the target P(Y) = `4/5`
Probability Z hitting the target P(Z) = `2/3`
`"P"(bar"X")` = 1 – P(X) = `1 - 3/4 = 1/4`
`"P"(bar"Y")` = 1 – P(Y) = `1 - 4/5 = 1/5`
`"P"(bar"Z")` = 1 – P(Z) = `1 - 2/3 = 1/3`
Probability hitting the target exactly by 2 hits
= `"P"[("X" ∩ "Y" ∩ bar"Z") ∪ (bar"X" ∩ "Y" ∩ "Z") ∪ ("X" ∩bar"Y" ∩ "Z")]`
= `"P"("X" ∩ "Y" ∩ bar"Z") + "P"(bar"X" ∩ "Y" ∩ "Z") + "P"("X" ∩bar"Y" ∩ "Z")`
= `"P"("X") "P"("Y") "P"(bar"Z") + "P"(bar"X") "P"("Y") "P"("Z") + "P"("X") "P"(bar"Y") "P"("Z")`
= `3/4 xx 4/5 xx 1/3 + 1/4 xx 4/5 xx 2/3 + 3/4 xx 1/5 xx 2/3`
= `(12 + 8 + 6)/60`
= `26/60`
= `13/30`
APPEARS IN
संबंधित प्रश्न
If P(A) = 0.8, P(B) = 0.5 and P(B|A) = 0.4, find P(A ∪ B)
Determine P(E|F).
A coin is tossed three times, where
E: at least two heads, F: at most two heads
Determine P(E|F).
A coin is tossed three times, where
E: at most two tails, F: at least one tail
A card is drawn from a well-shuffled pack of playing cards. What is the probability that it is either a spade or an ace or both?
Box I contains two white and three black balls. Box II contains four white and one black balls and box III contains three white ·and four black balls. A dice having three red, two yellow and one green face, is thrown to select the box. If red face turns up, we pick up the box I, if a yellow face turns up we pick up box II, otherwise, we pick up box III. Then, we draw a ball from the selected box. If the ball is drawn is white, what is the probability that the dice had turned up with a red face?
If events A and B are independent, such that `P(A)= 3/5`, `P(B)=2/3` 'find P(A ∪ B).
A bag contains 10 white balls and 15 black balls. Two balls are drawn in succession without replacement. What is the probability that, first is white and second is black?
Three fair coins are tossed. What is the probability of getting three heads given that at least two coins show heads?
Select the correct option from the given alternatives :
Bag I contains 3 red and 4 black balls while another Bag II contains 5 red and 6 black balls. One ball is drawn at random from one of the bags and it is found to be red. The probability that it was drawn from Bag II
Given P(A) = 0.4 and P(A ∪ B) = 0.7 Find P(B) if A and B are independent events
Given P(A) = 0.4 and P(A ∪ B) = 0.7 Find P(B) if P(B/A) = 0.5
Three machines E1, E2, E3 in a certain factory produced 50%, 25% and 25%, respectively, of the total daily output of electric tubes. It is known that 4% of the tubes produced one each of machines E1 and E2 are defective, and that 5% of those produced on E3 are defective. If one tube is picked up at random from a day’s production, calculate the probability that it is defective.
Find the probability that in 10 throws of a fair die a score which is a multiple of 3 will be obtained in at least 8 of the throws.
If P(A) = `4/5`, and P(A ∩ B) = `7/10`, then P(B|A) is equal to ______.
If P(A ∩ B) = `7/10` and P(B) = `17/20`, then P(A|B) equals ______.
Let A and B be two non-null events such that A ⊂ B. Then, which of the following statements is always correct?
If the sum of numbers obtained on throwing a pair of dice is 9, then the probability that number obtained on one of the dice is 4, is ______.
Which conditional probabilities are equal to 1 when \[P(F)\neq 0\]?
Which statement correctly describes the numerator in \[P(A\mid B)=\frac{P(A\cap B)}{P(B)}\]?
